OpenStudy (anonymous):

Help please(: A (4, 4), B (7, 0), C (11, 3), and D (8, 7) are four points on the coordinate grid. Miranda and Pete joined the points using straight lines to draw a quadrilateral ABCD. Miranda wrote the following statements to prove that, “ABCD is a parallelogram that is not a rhombus.” slope of AB=(4-0)/(4-70)=-4/3 slope of DC=(7-3)/(8-11)=-4/3 slope of BC=(0-3)/(7-11)=3/4 slope of AD=(4-7)/(4-8)=-3/4 Pete wrote the following statements to prove that, “ABCD is a rhombus.”

OpenStudy (anonymous):

one sec im typing the rest sorry its really long x.x

OpenStudy (anonymous):

(petes statements) AB=sqrt((4-7)^2+(4-0)^2=sqrt(25)=5 BC=sqrt((7-11)^2+(0-3)^2=sqrt(25)=5 CD=sqrt((11-8)^2+(3-7)^2=sqrt(25)=5 DA=sqrt((8-4)^2+(7-4)^2=sqrt(25)=5

OpenStudy (anonymous):

:( and im supposed to prove which statement is correct